# Multi-Modal Federated Learning for Cancer Staging over Non-IID Datasets with Unbalanced Modalities

### v1

@@ III-C3 Local ML Model Training
and $\bm{\omega}_{n}^{(t),k}=\big{\{}\{\bm{\omega}_{m,n}^{(t),k}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}^{(t),k}_{n}\big{\}}.$ In (6), $\mathcal{B}_{n}^{(t),k}$ denotes a mini-batch of data sampled uniformly at random from $\mathcal{D}_{n}^{\mathsf{Tr}}$ and $\eta^{(t),k}$ is the SGD learning-rate. Assuming that institutions conduct $K$ SGD iterations before transmitting their models to the server, we let $\bm{\omega}_{n}^{(t),K}=\big{\{}\{\bm{\omega}_{m,n}^{(t),K}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}^{(t),K}_{n}\big{\}}$ denote the final local encoder and decoder parameters at institution $n$.
@@ III-C4 Local Model Transfer and Global Model Aggregations
At the end of each local training round, after receiving the parameters of the encoders and decoders from the institutions, the server then aggregates them as follows to obtain a unique encoder parameter set for each modality and a unique decoder parameter set for each of the existing combinations of modalities in the system:
$\widehat{\bm{\omega}}_{m}^{(t+1)}$ $=\frac{\sum_{n=0}^{N}a_{m,n}|\mathcal{D}_{n}|\bm{\omega}_{m,n}^{(t),K}}{|\mathcal{D}|},\forall m\in\mathcal{M},$ (7)
$\widehat{\bm{\varpi}}_{C}^{(t+1)}$ $=\frac{\sum_{n=0}^{N}a_{C,n}|\mathcal{D}_{n}|\bm{\varpi}_{n}^{(t),K}}{|\mathcal{D}|},\forall C\in P(\mathcal{M}),$ (8)
where $a_{m,n}$, $\forall n,m$ is a binary indicator which takes the value of $1$ if institution $n$ has access to modality $m$; otherwise $a_{m,n}=0$. Similarly, $a_{C,n}$, $\forall n,C$, is a binary indicator, which is $1$ if institution $n$ has modality combination $C=\{m_{i}\}_{i\in\{1,\cdots,\mathcal{M}_{n}\}}$; otherwise $a_{C,n}=0$. After conducting the model aggregation step, the new sets of encoders and decoders are sent back in the downstream link to the institutions, which are then used to synchronize the local models. In particular, each institution $n$ with modality combination $C$ will conduct ${\small\bm{\omega}_{m,n}^{(t+1),0}\leftarrow\widehat{\bm{\omega}}_{m}^{(t+1)}}$ and ${\small\bm{\varpi}_{n}^{(t+1),0}\leftarrow\widehat{\bm{\varpi}}_{C}^{(t+1)}}$ (i.e., $\bm{\omega}_{n}^{(t+1),0}=\big{[}\{\widehat{\bm{\omega}}_{m}^{(t+1)}\}_{m\in\mathcal{M}_{n}}:\widehat{\bm{\varpi}}_{C}^{(t+1)}\big{]}$). This synchronization will be followed by SGD iterations described in (5) and then the next global aggregation step. This procedure continues until reaching a desired accuracy.
Fig. 3: Loss plots of the global models based on the modalities possessed across institutions in the first 100 global aggregation rounds. Plots consider various modality combinations: (a) image and mRNA modalities, (b) clinical and mRNA, and (c) image and clinical. The shaded areas represent the noise (20% standard deviation) for each combination of modalities tested. A key observation is the occasional degradation of the performance of the multi-modal ML models compared to their single-modal counterparts. This observation is consistent with the claims in [31] which emphasises on the necessity of syncing the convergence rates of the modalities in a multi-modal FL scenario. Further, our method leads to the balance of convergence rates across the modalities.
@@ III-D Key Challenges and Design Criteria
Our proposed system model contains nuances in decoupling the aggregation of encoders and decoders, which is necessary in handling institutions with an unbalanced number of modalities. However, achieving a high prediction performance across this system faces a major obstacle caused by straggling modalities. We demonstrate this phenomenon in Fig. 3, in which we conduct ML model training according to Sec. III-C, which implies training all institutions with unified learning-rates (labelled as Naïve FL in the plots). We consider a dataset of three modalities for cancer staging, consisting of mRNA sequences, histopathological images, and clinical information and presume the existence of $9$ institutions, where $3$ have access to only one modality of data, $3$ to two modalities of data, and $3$ to all modalities of data. The details of the dataset used for patients are given in Sec. V-A. As can be seen in the figure, the performance of some modalities is notably better than the others, as the losses are decreasing faster for those modalities but the same does not apply to others. Furthermore, during the local training phase, the ML model of some modality combinations has converged notably faster than the rest. Also, some modalities struggle to improve their loss and accuracy during the ML model training. We next aim to propose a framework to address these issues, the result of which is labelled as ours in the plots in Fig. 3. It can be observed from the curves that our method balances the convergence rates of various modality combinations, leading to a notable performance gain as verified later in Sec. V.
@@ IV-A Tackling Multi-Modal FL with Unbalanced Modalities
Considering the discussion in Sec. III-D, we aim to propose a methodology for multi-modal FL that alters the local gradient computations to solve the problem of different convergence rates of various modalities. We first introduce the two key components of our method:
  • Distributed Gradient Blending (DGB): To remove the barrier of convergence speeds, DGB weights the gradients that are being back-propagated through each encoder in an institution’s model to unify the convergence rate of different modalities. DGB achieves this through acquiring a set of train and validation losses from each institution at the end of each local ML training round, which is in turn used to calculate two metrics of overfitting and generalization. These two metrics are used to obtain the aforementioned weights for the gradients, which can accelerate or slacken the ML training rate of different modalities at various institutions.
  • Proximity-Aware Client Weighting (PCW): PCW mitigates the performance degradation of DGB caused by the heterogeneity of data across institutions, which leads to the bias of the locally trained ML models, especially as the length of local ML model training (i.e., $K$) grows. PCW tunes the weights/importance of ML models of institutions during the system-wide DGB weight calculation based on the trajectory of models’ –both local and global– gradients. Specifically, PCW assigns greater importance to institutions whose local ML parameters’ trajectory exhibit a closer alignment with that of the global model. This signifies that institutions with superior data quality (i.e., diverse local datasets that better encapsulate the global data distribution, and thus do not induce bias onto the locally trained ML model), will have a more pronounced influence on the calculation of DGB weights. By mitigating the impact of local ML model bias, PCW makes DGB robust against the increase in the number of local SGD iterations $K$ in each global aggregation (i.e., reductions in uplink transmission frequency).
@@ IV-B Distributed Gradient Blending (DGB)
We next derive two metrics called overfitting $O_{C}^{(t)}$ and generalization $G_{C}^{(t)}$ for each modality combination $C$ across the system, where $C\in\mathcal{P}(\mathcal{M})$, defined as
@@
where ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},{(t)}}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},{(t)}}$ are the average of the train and validation loss from institutions training on data with modality combination $C=\mathcal{M}_{n}$, obtained based on (IV-B) and (IV-B) as later formalized through PCW in Sec. IV-C. At each global aggregation step $t$, the above two metrics are used in defining another metric, called Distributed Overfitting to Generalization Ratio (DOGR) denoted by $\Gamma_{C}^{(t)}$, defined as follows:
@@
$\Gamma_{C}^{(t)}=\frac{1}{\varphi^{(t)}}\times\frac{\left|\Delta G_{C}^{(t)}\right|}{\left(\Delta O_{C}^{(t)}\right)^{2}},$ (13)
@@
where $\varphi^{(t)}=\sum_{C\in\mathcal{P}(\mathcal{M}_{n})}\frac{|\Delta G_{C}|}{2(\Delta O_{C})^{2}}$ is a normalization coefficient. Also, $\Delta O_{C}^{(t)}=O_{C}^{(t)}-O_{C}^{(t-1)}$ and $\Delta G_{C}^{(t)}=G_{C}^{(t)}-G_{C}^{(t-1)}$ represent the variations in overfitting and generalization, respectively. In (13), $\Delta O_{C}^{(t)}$ captures the change in the overfitting during the $t$-th global aggregation step for modality subset $C$ under the used step size. Also, $\Delta G_{C}^{(t)}$ captures the change in the validation loss during the latest global aggregation step for modality subset $C$, which measures the overall performance variation of the respective modality subset. Consequently, $\Gamma_{C}^{(t)}$ is a metric that takes smaller values when the generalization of the modality subset $C$ is decreasing or the model is becoming overfit during the global round and takes larger values otherwise. Thus, we can use $\Gamma_{C}^{(t)}$ to accelerate and slacken the learning-rate of different modality subsets according to their exhibited performance by increasing and dampening the unified SGD step-size during each local training round as discussed below.
@@
We incorporate $\Gamma_{C}^{(t)}$ in the learning-rate of the local SGD iterations at each institution $n$. In particular, we modify the local SGD iterations given by (5) to the following update rule, which accelerates/slackens the learning-rate across different modality subsets of various institutions ($10\leq k\leq K-$ 1):
$\bm{\omega}^{(t),k+1}_{m,n}=\bm{\omega}^{(t),k}_{m,n}-\eta^{(t),k}\Gamma_{m}^{(t)}\widetilde{\nabla}_{m}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n};\mathcal{B}_{n}^{(t),k}),\forall m\in\mathcal{M}_{n},$ (14)
@@
where ~$\widetilde{\nabla}_{m}$ is the stochastic gradient with respect to the parameters of modality $m$’s encoder (note $\Gamma_{m}^{(t)}=\Gamma_{C}^{(t)}$ with $C=m$). The classifier update is also modified as follows:
$\bm{\varpi}^{(t),k+1}_{n}=\bm{\varpi}^{(t),k}_{n}-\eta^{(t),k}\hskip 0.85358pt\Gamma_{\mathcal{M}_{n}}^{(t)}\widetilde{\nabla}_{\bm{\varpi}}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n};\mathcal{B}_{n}^{(t),k}),$ (15)
@@
where $\Gamma_{\mathcal{M}_{n}}^{(t)}$ ($\Gamma_{\mathcal{M}_{n}}^{(t)}=\Gamma_{C}^{(t)}$ with $C=\mathcal{M}_{n}$) captures the DOGR for the subset of modalities at institution $n$ and ~$\widetilde{\nabla}_{\bm{\varpi}}$ is the stochastic gradient with respect to its classifier parameters.
Naïve implementation of DGB for longer local round lengths (i.e., $K$) in our system will result in biased ML models (due to the heterogeneity of local datasets) being aggregated to estimate the gradient blending weights. Thus, one of the conditions that needs to be met in gradient blending is that the dataset used to estimate the weights should follow the same distribution as the general dataset that all of the institutions possess, which corresponds to the dataset $\mathcal{D}$ in our system. Meanwhile, one of the general phenomena in FL is the non-iid-ness of the local datasets. This, contrary to the requirements of DGB, makes the weights estimated unreliable in the sense that institutions with different ML model biases will have the same contribution to DGB. This is why we propose PCW in order to control the contribution of each institution’s loss to the gradient weight estimation based on the distance between their data distribution $\mathcal{D}_{n}$ and $\mathcal{D}$. In doing so, we will define a dynamic metric that conceptualizes the distance between institutions data distributions, and manipulates the weight with which each $\mathcal{L}_{n}^{\mathsf{Tr},(t)}$ and $\mathcal{L}_{n}^{\mathsf{Va},(t)}$ is taken into effect.
Remark 1 (Learning Methodology, Unbalanced Number of Modalities, Robustness to Non-iid data and Data Scarcity).
@@
The issue of varying convergence rates across modalities has also been observed in conventional centralized multi-modal ML [21]. To address this issue, [21] proposed a methodology, called gradient blending (GB), for centralized ML settings, which entails the use of a separate loss for each modality and a multi-modal model loss, weighting gradients non-uniformly when back-propagated over the ML model during training. As a result, since our methodology is inspired by [21], although it contains multiple nuances and differences compared to the method developed in [21], we call it Distributed Gradient Blending (DGB). In particular, our method differs significantly from the naïve extension of GB to the FL setting as done in [31]. In particular, existing methods [31, 21] use multiple sub-networks (i.e., one set of encoder and classifier for each modality), and thus, multiple losses to back-propagate during ML model training. In contrast, our work does not entail using multiple sub-networks; instead, we use only one classifier in each institution’s ML model alongside separate encoders dedicated to various modalities. We further use only one loss function and weigh its resulting gradient vector elements based on the encoder they belong to. This reduces the computation complexity of our method and the number of model parameters to train in each institution. Also, in addition to having a higher number of parameters, in the naïve implementations of gradient blending [21, 31], the methods require the train data available ($\mathcal{D}_{n}^{\mathsf{Tr}}$) to be split into two sections, one of which is used for the calculation of the overfitting to generalization ratio. This raises challenges when facing a scenario with scarce data, which is frequent in medical domains given the labor needed to annotate data. Our method, on the other hand, uses system-wide characteristics and values of the losses to calculate the DOGR which removes the necessity to have each of the train and validation data to be split into two at each institution, enabling more compatibility with data-scarce scenarios. More importantly, we address the side-effects of non-iid data on the performance of multi-modal FL through PCW, which makes DGB outperform the existing methodologies.

### v2

@@ III-C3 Local ML Model Training
and $\bm{\omega}_{n}^{(t),k}=\big{\{}\{\bm{\omega}_{m,n}^{(t),k}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}^{(t),k}_{n}\big{\}}.$ In (6), $\mathcal{B}_{n}^{(t),k}$ denotes a mini-batch of data sampled uniformly at random from $\mathcal{D}_{n}^{\mathsf{Tr}}$ and $\eta^{(t),k}$ is the learning-rate. Assuming that institutions conduct $K$ SGD iterations before transmitting their models to the server, we let $\bm{\omega}_{n}^{(t),K}=\big{\{}\{\bm{\omega}_{m,n}^{(t),K}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}^{(t),K}_{n}\big{\}}$ denote the final local encoder and classifier parameters at institution $n$.
@@ III-C4 Local Model Transfer and Global Model Aggregations
At the end of each local training round, after receiving the parameters of the encoders and classifiers from the institutions, the server then aggregates them as follows to obtain a unique encoder parameter set for each modality and a unique classifier parameter set for each of the existing combinations of modalities in the system:
$\widehat{\bm{\omega}}_{m}^{(t+1)}$ $=\frac{\sum_{n\in\mathcal{N}}a_{m,n}|\mathcal{D}_{n}|\bm{\omega}_{m,n}^{(t),K}}{\sum_{n\in\mathcal{N}}a_{m,n}|\mathcal{D}_{n}|},\forall m\in\mathcal{M},$ (7)
$\widehat{\bm{\varpi}}_{C}^{(t+1)}$ $=\frac{\sum_{n\in\mathcal{N}}b_{C,n}|\mathcal{D}_{n}|\bm{\varpi}_{n}^{(t),K}}{\sum_{n\in\mathcal{N}}b_{C,n}|\mathcal{D}_{n}|},\forall C\in\mathcal{P}(\mathcal{M}),$ (8)
where $a_{m,n}$, $\forall n,m$ is a binary indicator which takes the value of $1$ if institution $n$ has access to modality $m$; otherwise $a_{m,n}=0$. Similarly, $b_{C,n}$ is a binary indicator that takes the value of $1$ when the modalities at institution $n$ match the combination $C$ (i.e., $\mathcal{M}_{n}=C$); otherwise $b_{C,n}=0$. After conducting the model aggregation step, the new sets of encoders and classifiers are sent back in the downstream link to the institutions, which are then used to synchronize the local models. In particular, each institution $n$ with modality combination $C$ will conduct ${\small\bm{\omega}_{m,n}^{(t+1),0}\leftarrow\widehat{\bm{\omega}}_{m}^{(t+1)}}$ and ${\small\bm{\varpi}_{n}^{(t+1),0}\leftarrow\widehat{\bm{\varpi}}_{C}^{(t+1)}}$ (i.e., $\bm{\omega}_{n}^{(t+1),0}=\big{[}\{\widehat{\bm{\omega}}_{m}^{(t+1)}\}_{m\in\mathcal{M}_{n}}:\widehat{\bm{\varpi}}_{C}^{(t+1)}\big{]}$). This synchronization will be followed by SGD iterations described in (5) and then the next global aggregation step. This procedure continues until reaching a desired accuracy.
Fig. 3: Loss plots of the global models based on the modalities possessed across institutions in the first 100 global aggregation steps. Plots consider various modality combinations: (a) image and mRNA modalities, (b) clinical and mRNA, and (c) image and clinical. The shaded areas represent the noise (20% standard deviation) for each combination of modalities tested. A key observation is the occasional degradation of the performance of the multi-modal ML models compared to their single-modal counterparts. This observation is consistent with the claims in [31] which emphasises on the necessity of syncing the convergence rates of the modalities in a multi-modal FL scenario. Further, our method leads to the balance of convergence rates across the modalities.
@@ III-D Key Challenges and Design Criteria
Our proposed system model contains nuances in decoupling the aggregation of encoders and classifiers, which is necessary in handling institutions with an unbalanced number of modalities. However, achieving a high prediction performance across this system faces a major obstacle caused by straggling modalities. We demonstrate this phenomenon in Fig. 3, in which we conduct ML model training according to Sec. III-C, which implies training all institutions with unified learning-rates (labelled as Naïve FL in the plots). We consider a dataset of three modalities for cancer staging, consisting of mRNA sequences, histopathological images, and clinical information and presume the existence of $9$ institutions, where $3$ have access to only one modality of data, $3$ to two modalities of data, and $3$ to all modalities of data. The details of the dataset used for patients are given in Sec. V-A. As can be seen in the figure, the performance of some modalities is notably better than the others, as the losses are decreasing faster for those modalities but the same does not apply to others. Furthermore, during the local training phase, the ML model of some modality combinations has converged notably faster than the rest. Also, some modalities struggle to improve their loss and accuracy during the ML model training. We next aim to propose a framework to address these issues, the result of which is labelled as ours in the plots in Fig. 3. It can be observed from the curves that our method balances the convergence rates of various modality combinations, leading to a notable performance gain as verified later in Sec. V.
@@ IV-A Tackling Multi-Modal FL with Unbalanced Modalities
Considering the discussion in Sec. III-D, we aim to propose a methodology for multi-modal FL that alters the local gradient computations to solve the problem of different convergence rates of various modalities. In a nutshell, our method consists of two components: (i) distributed gradient blending (DGB), which is a methodology to evaluate the contribution of various modalities to the global model training without transferring the data across the network, which is then used to control the convergence rates of modalities so as to accelerate the FL model convergence; and (ii) proximity-aware client weighting (PCW), which is a method for tackling the impact of non-iid data and biased local datasets at the institutions via evaluating the quality of data possessed by institutions – without data transfer across the network – via a similarity metric. This metric is then used to quantify the extent of reliance on the performance of each local model on calculating the DGB parameters. We next provide a high-level description of these two key components of our method, and mathematically formalize them later in Sec. IV-B and IV-C:
  • Distributed Gradient Blending (DGB): To remove the barrier of non-uniform convergence speeds across various data modalities, DGB weighs the gradients that are being back-propagated through each encoder in an institution’s model to unify the convergence rates of modalities. DGB achieves this through acquiring a set of train and validation losses from each institution at the end of each local ML training round, which is in turn used to calculate two metrics of overfitting and generalization. These two metrics are then used to calculate the distributed overfitting to generalization ratio (DOGR), formalized in Sec. IV-B. The DOGR will measure the change in overfitting and generalization ability of the present modalities across the available institutions to quantify the average performance improvement for each modality combination across institutions in the recent global aggregation step. DOGR is then integrated to obtain the aforementioned weights for the gradients, which can accelerate or slacken the ML training rate of different modalities at various institutions.
  • Proximity-Aware Client Weighting (PCW): PCW mitigates the performance degradation of DGB caused by the heterogeneity of data across institutions, which leads to the bias of the locally trained ML models, especially as the length of local ML model training (i.e., $K$) grows. To overcome the effect of non-iid data distribution across institutions on the losses acquired from them during the calculation of DOGR weights, PCW aims to quantify the extent of reliance on the local losses of various institutions – used in computing the DOGR parameters – according to the quality of data held by institutions. To this end, PCW weighs the received local losses of institutions at the server based on the trajectory of both local and global models. Specifically, PCW promotes more reliance on the local losses of institutions whose local ML parameters’ trajectory exhibits a closer alignment with that of the global model. This signifies that institutions with superior data quality (i.e., diverse local datasets that better encapsulate the global data distribution, and thus do not induce bias onto the locally trained ML model), will have a more pronounced influence on the calculation of DGB weights. By mitigating the impact of local ML model bias, PCW makes DGB robust against the increase in the number of local SGD iterations $K$ in each global aggregation (i.e., reductions in uplink transmission frequency). The mathematical explanation of PCW is discussed in Sec. IV-C.
@@ IV-B Distributed Gradient Blending (DGB)
We next derive two metrics called overfitting $O_{C}^{(t)}$ and generalization $G_{C}^{(t)}$ for each modality combination $C$ across the system, where $C\in\mathcal{P}(\mathcal{M})$, defined as [^7] In the following, we assume that at least one institution with single-modal data for each modality $m$ (i.e., institution $n$ with $\mathcal{M}_{n}=\{m\}$) exists in the system. In other words, single-modal institutions cover all the modalities of data present in the system.
@@
where ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},{(t)}}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},{(t)}}$ are the average/aggregation of the train and validation loss from institutions training on data with modality combination $C$ (i.e., from institutions $n$ such that $\mathcal{M}_{n}=C$), obtained based on (IV-B) and (IV-B) as later formalized through PCW in Sec. IV-C. It is worth mentioning that in the previous approaches toward gradient blending, e.g., OGR metric in [21], these metrics were calculated locally for each institution, where the losses in (IV-B) and (IV-B) were directly used in (11) and (12) without averaging. [^8] This will make the methodology prone to the bias of local datasets in our setting. In particular, biased and scarce local data sets at various institutions, which are common in FL and medical settings, will render localized computation of these parameters followed by their local adaptation for tuning the learning-rates of various modalities ineffective. For example, one modality may contribute to the training highly in one institution and also have a higher loss, making the trained model sensitive to the features of that specific modality. However, on a global scale, that modality may not have as high contribution and may also possess a small loss value when considered all the institutions.
After obtaining $O_{C}^{(t)}$ and $G_{C}^{(t)}$ for all $C\in\mathcal{P}(\mathcal{M})$ at the server and broadcasting them across the institutions, at each global aggregation step $t$, these metrics are used in defining another metric, called Distributed Overfitting to Generalization Ratio (DOGR) for each institution $n$. In particular, each institution $n$ only uses a subset of $O_{C}^{(t)}$ and $G_{C}^{(t)}$ received from the server: only the values corresponding to the local encoders (i.e., $O_{m}^{(t)}$ for all $m\in\mathcal{M}_{n}$ obtained via $O_{m}^{(t)}=O_{C}^{(t)}$ when $C=\{m\}$; $G_{m}^{(t)}$ for all $m\in\mathcal{M}_{n}$ obtained via $G_{m}^{(t)}=G_{C}^{(t)}$ when $C=\{m\}$) and the overall combination held by the institution (i.e., $O_{\mathcal{M}_{n}}^{(t)}$ obtained via $O_{\mathcal{M}_{n}}^{(t)}=O_{C}^{(t)}$ when $C=\mathcal{M}_{n}$; $G_{\mathcal{M}_{n}}^{(t)}$ obtained via $G_{\mathcal{M}_{n}}^{(t)}=G_{C}^{(t)}$ when $C=\mathcal{M}_{n}$). Subsequently, the DOGR at each institution $n$ is computed for each modality $m$ (i.e., corresponding to each encoder) and the overall modality set held by the institution $\mathcal{M}_{n}$ (i.e., corresponding to the classifier) as follows:
@@
$\Gamma_{m,n}^{(t)}=\frac{1}{\varphi^{(t)}_{n}}\times\frac{\left|\Delta G_{m}^{(t)}\right|^{2}}{\left(\Delta O_{m}^{(t)}\right)^{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ m\in\mathcal{M}_{n},$ (13)
$\Gamma_{\mathcal{M}_{n},n}^{(t)}=\frac{1}{\varphi^{(t)}_{n}}\times\frac{\left|\Delta G_{\mathcal{M}_{n}}^{(t)}\right|^{2}}{\left(\Delta O_{\mathcal{M}_{n}}^{(t)}\right)^{2}}.$
@@
where $\varphi_{n}^{(t)}=\frac{|\Delta G_{\mathcal{M}_{n}}|^{2}}{2(\Delta O_{\mathcal{M}_{n}})^{2}}+\sum_{m\in\mathcal{M}_{n}}\frac{|\Delta G_{m}|^{2}}{2(\Delta O_{m})^{2}}$ is a normalization coefficient obtained over the encoder modalities and the classifier combination existing in institution $n$. Also, $\Delta O_{m}^{(t)}=O_{m}^{(t)}-O_{m}^{(t-1)}$ and $\Delta G_{m}^{(t)}=G_{m}^{(t)}-G_{m}^{(t-1)}$, where $m\in\mathcal{M}_{n}$, represent the variations in overfitting and generalization, respectively (similarly $\Delta O_{\mathcal{M}_{n}}^{(t)}=O_{\mathcal{M}_{n}}^{(t)}-O_{\mathcal{M}_{n}}^{(t-1)}$ and $\Delta G_{\mathcal{M}_{n}}^{(t)}=G_{\mathcal{M}_{n}}^{(t)}-G_{\mathcal{M}_{n}}^{(t-1)}$). In particular, $\Delta O_{m}^{(t)}$ captures the change in the overfitting during the $t$-th global aggregation step for modality $m$ under the used step size. Also, $\Delta G_{m}^{(t)}$ captures the change in the validation loss during the latest global aggregation step for modality $m$, which measures the overall performance variation of the respective modality. Consequently, $\Gamma_{m,n}^{(t)}$ is a metric that takes smaller values when the generalization of the modality encoder $m$ present at institution $n$ is decreasing or the model is becoming overfit during the global round and takes larger values otherwise. Thus, we can use $\Gamma_{m,n}^{(t)}$ to accelerate and slacken the learning-rate of different modality encoders present in institution $n$ according to their exhibited performance by increasing and dampening the unified SGD step-size during each local training round as discussed below. The same discussion holds in the case of $\Delta O_{\mathcal{M}_{n}}^{(t)}$, $\Delta G_{\mathcal{M}_{n}}^{(t)}$, and $\Gamma_{\mathcal{M}_{n},n}^{(t)}$ which are used to increase/decrease the SGD step size in the classifier of each institution $n$.
@@
We incorporate DOGR in the learning-rate of the local SGD iterations at each institution $n$. In particular, we modify the local SGD iterations given by (5) to the following update rule, which accelerates/slackens the learning-rate across different modality subsets of various institutions ($10\leq k\leq K-$ 1):
$\bm{\omega}^{(t),k+1}_{m,n}=\bm{\omega}^{(t),k}_{m,n}-\eta^{(t),k}\Gamma_{m,n}^{(t)}\widetilde{\nabla}_{m}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n};\mathcal{B}_{n}^{(t),k}),\forall m\in\mathcal{M}_{n},$ (14)
@@
where ~$\widetilde{\nabla}_{m}$ is the stochastic gradient with respect to the parameters of modality $m$’s encoder. The classifier update is also modified as
$\bm{\varpi}^{(t),k+1}_{n}=\bm{\varpi}^{(t),k}_{n}-\eta^{(t),k}\hskip 0.85358pt\Gamma_{\mathcal{M}_{n},n}^{(t)}\widetilde{\nabla}_{\bm{\varpi}}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n};\mathcal{B}_{n}^{(t),k}),$ (15)
@@
where ~$\widetilde{\nabla}_{\bm{\varpi}}$ is the stochastic gradient with respect to the classifier parameters.
To make DGB applicable for FL scenarios it should be adapted to one of the most unique aspects of FL settings, which is the existence of non-iid and biased local datasets across the institutions. [^9] This further results in local ML bias specifically when longer local round lengths $K$ are used. In particular, obtaining (11) and (12) with naïve averaging of the local loss values of various modality combinations in the system renders the computed values ineffective under non-iid data. For example, consider an institution where the distribution of data can be biased (e.g., containing disproportionately high ratio of data from specific classes/labels, or containing disproportionately high ratio of data from specific patient demographics). This will make the local ML model of that specific institution overfit to the biased local data, which will result in a noticeably low local validation loss, severely impacting the aggregated losses used in DOGR upon naïve averaging of the loss values across institutions to obtain (11) and (12). As a result, we have to migrate from conventional averaging of the losses across all institutions to the methods that can capture the quality of local datasets across institutions when aggregating their loss values. We achieve this through introducing PCW, which controls the contribution of each institution’s loss to the DGB parameter estimation based on the distance between its local data distribution $\mathcal{D}_{n}$ and the global data distribution $\mathcal{D}$. In doing so, in PCW, we will define a dynamic metric that conceptualizes the distance between institutions data distributions, and manipulates the impact with which each $\mathcal{L}_{n}^{\mathsf{Tr},(t)}$ and $\mathcal{L}_{n}^{\mathsf{Va},(t)}$ is taken into effect.
Remark 1.
@@
[Learning Methodology, Unbalanced Number of Modalities, Robustness to Non-iid data and Data Scarcity] The issue of varying convergence rates across modalities has also been observed in conventional centralized multi-modal ML [21]. To address this issue, [21] proposed a methodology, called gradient blending (GB), for centralized ML settings, which entails the use of a separate loss for each modality and a multi-modal model loss, weighting gradients non-uniformly when back-propagated over the ML model during training. As a result, since our methodology is inspired by [21], although it contains multiple nuances and differences compared to the method developed in [21], we call it Distributed Gradient Blending (DGB). In particular, our method differs significantly from the naïve extension of GB to the FL setting as done in [31]. In particular, existing methods [31, 21] use multiple sub-networks (i.e., one set of encoder and classifier for each modality), and thus, multiple losses to back-propagate during ML model training. In contrast, our work does not entail using multiple sub-networks; instead, we use only one classifier in each institution’s ML model alongside separate encoders dedicated to various modalities. We further use only one loss function and weigh its resulting gradient vector elements based on the encoder they belong to. This reduces the computation complexity of our method and the number of model parameters to train in each institution. Also, in addition to having a higher number of parameters, in the naïve implementations of gradient blending [21, 31], the methods require the train data available ($\mathcal{D}_{n}^{\mathsf{Tr}}$) to be split into two sections, one of which is used for the calculation of the overfitting to generalization ratio. This raises challenges when facing a scenario with scarce data, which is frequent in medical domains given the labor needed to annotate data. Our method, on the other hand, uses system-wide characteristics and values of the losses to calculate the DOGR which removes the necessity to have each of the train and validation data to be split into two at each institution, enabling more compatibility with data-scarce scenarios. This is specifically shown in our method when using the averaged loss of the local ML models for each combination, instead of using the local loss for each local model separately. More importantly, we address the side-effects of non-iid data on the performance of multi-modal FL through PCW which is a new condition that arises because of the new configuration of DGB, aiming to elevate its performance.